# Need for zkSpeed: Accelerating HyperPlonk for Zero-Knowledge Proofs

### v1

@@ Acknowledgements.
- hbm (2023) 2023. High Bandwidth Memory DRAM (HBM3). Technical Report JESD238A. JEDEC. [Online]. Available: https://www.jedec.org/standards-documents/docs/jesd238a.
@@
- Morgan (2019b) Timothy Prickett Morgan. 2019b. A Deep Dive Into AMD’s Rome EPYC Architecture. https://www.nextplatform.com/2019/08/15/a-deep-dive-into-amds-rome-epyc-architecture/ Accessed: [Insert date of access here].
@@
- Pornin (2020) Thomas Pornin. 2020. Optimized Binary GCD for Modular Inversion. 2020/972 (2020). https://eprint.iacr.org/2020/972 Publication info: Preprint. MINOR revision..
@@
- Yang et al. (2025) Zhengbang Yang, Lutan Zhao, Peinan Li, Han Liu, Kai Li, Boyan Zhao, Dan Meng, and Rui Hou. 2025. LegoZK: a Dynamically Reconfigurable Accelerator for Zero Knowledge Proof. (2025).
- Zhang et al. (2021) Ye Zhang, Shuo Wang, Xian Zhang, Jiangbin Dong, Xingzhong Mao, Fan Long, Cong Wang, Dong Zhou, Mingyu Gao, , and Guangyu Sun. 2021. PipeZK: Accelerating Zero-Knowledge Proof with a Pipelined Architecture. In 2021 ACM/IEEE 48th Annual International Symposium on Computer Architecture (ISCA).
- Zhang et al. ([n. d.]) Zhenfei Zhang, Binyi Chen, Benedikt Bünz, and Alex Xiong. [n. d.]. Hyperplonk Implementation.

### v2

@@ Acknowledgements.
- hbm (2023) 2023. High Bandwidth Memory DRAM (HBM3). Technical Report JESD238A. JEDEC. https://www.jedec.org/standards-documents/docs/jesd238a.
@@
- Abbaszadeh et al. (2024) Kasra Abbaszadeh, Christodoulos Pappas, Jonathan Katz, and Dimitrios Papadopoulos. 2024. Zero-Knowledge Proofs of Training for Deep Neural Networks. In Proceedings of the 2024 on ACM SIGSAC Conference on Computer and Communications Security (Salt Lake City, UT, USA) (CCS ’24). Association for Computing Machinery, New York, NY, USA, 4316–4330. https://doi.org/10.1145/3658644.3670316
@@
- Chillotti et al. (2020) Ilaria Chillotti, Nicolas Gama, Mariya Georgieva, and Malika Izabachène. 2020. TFHE: fast fully homomorphic encryption over the torus. Journal of Cryptology 33, 1 (2020), 34–91.
@@
- Ebel et al. (2025) Austin Ebel, Karthik Garimella, and Brandon Reagen. 2025. Orion: A Fully Homomorphic Encryption Framework for Deep Learning. In Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (Rotterdam, Netherlands) (ASPLOS ’25). Association for Computing Machinery, New York, NY, USA, 734–749. https://doi.org/10.1145/3676641.3716008
@@
- Feng et al. (2024) Boyuan Feng, Zheng Wang, Yuke Wang, Shu Yang, and Yufei Ding. 2024. ZENO: A Type-based Optimization Framework for Zero Knowledge Neural Network Inference. In Proceedings of the 29th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1 (La Jolla, CA, USA) (ASPLOS ’24). Association for Computing Machinery, New York, NY, USA, 450–464. https://doi.org/10.1145/3617232.3624852
@@
- Jha and Reagen (2024) Nandan Kumar Jha and Brandon Reagen. 2024. DeepReShape: Redesigning Neural Networks for Efficient Private Inference. Transactions on Machine Learning Research (TMLR) (2024).
@@
- Mo et al. (2025a) Jianqiao Mo, Alhad Daftardar, Joey Ah-kiow, Kaiyue Guo, Benedikt Bünz, Siddharth Garg, and Brandon Reagen. 2025a. MTU: The Multifunction Tree Unit in zkSpeed for Accelerating HyperPlonk. arXiv:2507.16793 [cs.AR] https://arxiv.org/abs/2507.16793
@@
- Mo et al. (2025b) Jianqiao Cambridge Mo, Karthik Garimella, Austin Ebel, and Brandon Reagen. 2025b. ABLE: Optimizing Mixed Arithmetic and Boolean Garbled Circuit. Cryptology ePrint Archive, Paper 2025/048. https://eprint.iacr.org/2025/048
- Monnot (2024) Barnabé Monnot. 2024. On Block Sizes, Gas Limits, and Scalability. https://ethresear.ch/t/on-block-sizes-gas-limits-and-scalability/18444. Accessed: 2025-05-06.
@@
- Morgan (2019b) Timothy Prickett Morgan. 2019b. A Deep Dive Into AMD’s Rome EPYC Architecture. https://www.nextplatform.com/2019/08/15/a-deep-dive-into-amds-rome-epyc-architecture/
@@
- Pornin (2020) Thomas Pornin. 2020. Optimized Binary GCD for Modular Inversion. 2020/972 (2020). https://eprint.iacr.org/2020/972
@@
- Yang et al. (2025) Zhengbang Yang, Lutan Zhao, Peinan Li, Han Liu, Kai Li, Boyan Zhao, Dan Meng, and Rui Hou. 2025. LegoZK: A Dynamically Reconfigurable Accelerator for Zero-Knowledge Proof. In 2025 IEEE International Symposium on High Performance Computer Architecture (HPCA). 113–126. https://doi.org/10.1109/HPCA61900.2025.00020
- Zhang and Franchetti (2025) Naifeng Zhang and Franz Franchetti. 2025. Code Generation for Cryptographic Kernels using Multi-word Modular Arithmetic on GPU. In Proceedings of the 23rd ACM/IEEE International Symposium on Code Generation and Optimization (Las Vegas, NV, USA) (CGO ’25). Association for Computing Machinery, New York, NY, USA, 476–492. https://doi.org/10.1145/3696443.3708948
- Zhang et al. (2021) Ye Zhang, Shuo Wang, Xian Zhang, Jiangbin Dong, Xingzhong Mao, Fan Long, Cong Wang, Dong Zhou, Mingyu Gao, and Guangyu Sun. 2021. PipeZK: Accelerating Zero-Knowledge Proof with a Pipelined Architecture. In 2021 ACM/IEEE 48th Annual International Symposium on Computer Architecture (ISCA).
- Zhang et al. (2022) Zhenfei Zhang, Binyi Chen, Benedikt Bünz, and Alex Xiong. 2022. Hyperplonk Implementation. https://github.com/EspressoSystems/hyperplonk